OpenStudy (anonymous):

solve using the perfect square factoring patterns: y^2 + 18y + 81 = 0

OpenStudy (nurali):

y^2 +18y +81 = 0 Again rudimentary knowledge of time tables should tell you we're dealing with the number 9. 9+9 = 18, 9*9 = 81 The equation says +18y and +81 so clearly the sign has to be positive. (y+9)^2 = 0 (y+9) = square root of 0 Ok this is going to break with normal quadratic equations as it will only have one solution, when we square root zero, we get +ve 0 and -ve 0 but they both mean zero. (y+9) = 0 y= -9

OpenStudy (anonymous):

answers a)9 b)9, -9 c)-9 d)0, -1

OpenStudy (anonymous):

ok thank you

OpenStudy (nurali):

Anytime